OpenStudy (anonymous):

Write an equivalent form of 6 as a fifth root.

OpenStudy (kinggeorge):

\[6=\sqrt[5]{x}\] so \[x = 6^5 = 7776\] so \[6 = \sqrt[5]{7776}\]

OpenStudy (anonymous):

thanks (: